Powell Products manufactures its products in two separate departments: Machining and Assembly. Total manufacturing overhead cRequirement 4. Compute the total amount of overhead allocated to each job if the company uses its current plantwide overheadRequirement 1. Compute the total manufacturing cost and sales price current plantwide Enter the percentage as a whole num JobBased on the departmental overhead rates and the sales price determined in Requirement 7, how much profit did the company reaMachining Assembly Department Department Job 500 4 MH 10 DL hours 2 DL hours Job 501 ... 8 10 DL hours 2 DL hours Both Jobs 5please ignore the answers I have put so far, I think they are wrong
Answer #1

1.Predetermined Plantwide rate = Total Overheads/Total labor hours

=1050000/10,000 = $105 per labor hour

2.Rate = Departmental Overheads/Driver

Machining = 650,000/5000 = $130 per machine hour

Department B = 400,000/8000 = $50 per labor hour

501 uses more

4 more machine hours, costs the company more resources

Job 500

Job 501

Total direct labor hours

12

12

Plantwide

     105.00

105.00

Overhead

1,260

1,260

Job 500

Job 501

Machining

=130*4 = $520

1,040

Assembly

500

500

Total

1,020

1,540

Equal overheads

More to 501

extra machine hours used

Correct/fair

Job 500

Job 501

Direct Material

1000

1000

Direct labor

240

240

Manufacturing overhead

1,260

1,260

Total manufacturing costs

2500

2500

Mark up

110%

110%

Sales Price

2,750

2,750

Job 500

Job 501

Sales Price

2750

2750

Total manufacturing costs

2500

2500

Gross Profit

250

250

Job 500

Job 501

Sales Price

2750

2750

Less: Costs

Direct Material

1000

1000

Direct labor

240

240

Manufacturing overhead

1,020

2260

1,540

2,780

Gross Profit

490

-30

Believes that it earns

Realises that loses money on 501